OpenStudy (anonymous):

First answer will get medal!! What happens when you multiply tan and cos, cos and sin, sec and sin.

OpenStudy (anonymous):

tan and cos: tan=sin/cos sin*cos/cos= sin/1. cos and sin: cos*sin=cos*sin, no simplification here. sec and sin: sec=1/cos, sin*(1/cos)=sin/cos=tan.

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

i dont get the tan and cos

OpenStudy (anonymous):

so tan = sin/cos. (sin/cos) * cos = (sin*cos)/(cos). because there is a cos in the numerator and denominator, you can cross the cos's off, leaving you with sin/1.

OpenStudy (anonymous):

why does tan equal sin / cos

OpenStudy (anonymous):

that's just the definition of tan

OpenStudy (anonymous):

sec= 1/cos csc= 1/sin tan= sin/cos cot= cos/sin sin is normal, just sin cos is normal, just cos all of them are related to sin and cos

OpenStudy (anonymous):

ok